OpenStudy (anonymous):

simplified version of 1/2 logsub 3 x + 2 log sub 3 y + 4 logsub 3 z

OpenStudy (anonymous):

Did you mean this? \[(1/2)*\log_3(x) + (2)*\log_3(y) + (4)*\log_3(z) \]

OpenStudy (anonymous):

There are 3 different ways to simplify that I can think of, would you like the simplest denominator?

OpenStudy (anonymous):

no. i mean 1/2 log sub 3 x+2 log sub3 y+4 log sub3 z and sure

OpenStudy (anonymous):

?

OpenStudy (anonymous):

In base ten (decimal # system), it simplifies to: \[( \log(x) + 4\log(y) + 8\log(z) ) / \log(9)\] A helpful rule for logarithms that is relevant here: \[\log_A X = (\log X) / (logA)\] For example: \[\log_32 = \log 2 / \log 3\] The default if there's no base specified is log base 10. Does that help?

OpenStudy (anonymous):

the answer should have log base 3 = the sqrt of x something

OpenStudy (anonymous):

?

OpenStudy (anonymous):

this is the problem ((1)/(2))log[3]x+2log[3]y+4log[3]z

OpenStudy (anonymous):

hereq

OpenStudy (anonymous):

do u know that a * log (x) = log (x^a)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

example 5 * log 3 = log (3^5)

OpenStudy (anonymous):

Is this listed an equation you're solving for? I was under the impression it was an expression you needed simplified... Based on what you initially wrote as your question :-/

OpenStudy (anonymous):

ok my answer choices are

OpenStudy (anonymous):

If it's like this: \[\log_3 x = ((1)/(2))\log_3 x+2 \log_3 y +4 \log_3 z\] This can be solved in terms of log base 3

OpenStudy (anonymous):

log base 3 sqrt x/ y^2z^4 log base 3 sqrt x y^2/ z^4 log base 3 sqrt x z^4/y^2 log base 3 sqrt x y^2 z^4

OpenStudy (anonymous):

4th one

OpenStudy (anonymous):

why? @agentx5 is this correct

OpenStudy (anonymous):

This will be kind of messy, but these following equations are identical. For base 3: \[(y^2 z^4 \log(\sqrt(x)))/(\log(3)) = (\log(x))/(2 \log(3))+(2 \log(y))/(\log(3))+(4 \log(z))/(\log(3))\] For base 10: \[y^2 z^4 \log(x) = \log(x)+4 \log(y)+8 \log(z)\] Read what I wrote above with change of base formula, it helps. Choice #4/D seems the most correct, based on what I now understand about your question. I concur. A helpful tip: when dealing with logarithms in particular, it's wise for one to be extra careful with parenthesis Does this answer your question sufficiently I hope? :D

OpenStudy (anonymous):

as much as a possible!